# S4ST: A Strong, Self-transferable, faSt, and Simple Scale Transformation for Data-free Transferable Targeted Attack

### v2

@@ 2.3 Input Transformations for Transferability
2.4 Failures of Existing Consensuses
We scrutinize the prevailing consensuses commonly referenced in the development of transformation methods for transfer attacks. Our objective is to ascertain their robustness as metrics reflecting targeted transferability and to determine their utility in guiding the design of transformations for TTAs. To achieve this, we assess the black-box transferability of various transformation techniques under both default and expanded hyperparameter configurations. Here, the black-box transferability is calculated by the variance in predicted probabilities of black-box models with respect to the target class $y_{t}$ between the adversarial (generated by $\mathcal{T}$-TMI) and clean samples:
We then examine the presence of a positive (or linear) correlation between their transferability and the quantitative measures derived from the established consensus. The consensuses under investigation and their corresponding measures are as follows:
  • Diversity. Initially, researchers emphasized the need for transformations that preserve loss [45, 10, 44, 9], ensuring that the surrogate model’s output remains relatively constant post-transformation. The intensity of transformations was thus limited to a narrow scope to maintain this characteristic. However, more recent advancements have shifted towards utilizing gradient averaging over multiple transformed samples (20$\times$) to achieve stable gradient updates [12, 11, 13, 46]. This shift implies that increased diversity might lead to improved transferability. We quantify diversity by calculating the average cross-entropy loss values of the surrogate model across 50,000 images from the ImageNet-1K validation set.
  • Attention Deviation. Dong et al. first pointed out that discrepancies in discriminative regions between models and defensive strategies can hinder adversarial transferability [9]. They advocated for the use of translation transformations to extend the impact of adversarial perturbations to wider attention regions. Following this, other studies have also utilized attention as a universal objective for crafting transferable attacks [47, 48]. More recently, BSR [12] and H-Aug [27] have set the new SOTAs for untargeted and targeted attacks by disrupting attention through image block shuffling and random attention-deviation transformations, respectively. We adopt the attention deviation metric proposed by Wei et al. [27], which computes the Intersection over Union (IoU) of Grad-CAM [49] activations concerning the target label between original and transformed images. The results are averaged over 1,000 images with distinct target labels.
The results are presented in Fig. 3. Within the intra-method analyses—where the relationship between different hyperparameter configurations of individual methods is evaluated—a positive or linear relationship may be discernible. Notably, the expanded parameters for the DI method expose the constraints of experimental setups in previous research. Nevertheless, this only suggests that these metrics can provide empirical guidance for parameter scaling within specific transformations. Importantly, there is a significant confusion of these metrics across different methods, which complicates the task of developing new transformation methods from the ground up.
@@ 3 The Proposed Method
This section systematically delineates the proposed method. It integrates the motivation derived from previous analyses, delineates two plausible strategies for gauging the impact of transformation methods on targeted transferability, and culminates in the introduction of a powerful yet efficient transformation technique to augment data-free TTAs.
@@ 3.1 Motivation
In our research, we fervently promote the investigation into robust blind estimators for transformation-induced targeted transferability. To this end, we aspire to establish design principles for formulating transformation techniques. Our inquiry specifically targets questions such as which transformations prove most effective and how to determine the synergistic effects of combining various transformations. We anticipate that our findings will identify transformations that facilitate highly effective data-free TTAs. Put succinctly, our goal is to surpass the efficacy of data-reliant methods while preserving efficiency and the ability to handle arbitrary (sample, label) pairs. This is accomplished without resorting to additional data or retraining models, thereby striving for an optimal balance between performance and practicality. Moreover, the estimators in question should conform to the prerequisites of black-box transfer attacks, thereby ensuring their applicability to a broader spectrum of data beyond natural imagery. To begin, we focus on basic transformations and subsequently investigate their potential synergistic enhancements. The specifics of the basic transformations evaluated and their respective intensities $\mathcal{S}$ are meticulously defined and expounded upon in Appendix A.
3.2 Assessing Effectiveness via Surrogate Self-Alignment
First, we manage to assess the effectiveness of attacking basic transformations to enhance the targeted transferability. This investigation is challenging due to the prohibition of direct evaluations on black-box models. To circumvent this, we suggest an approach rooted in the concept of model alignment. The underlying rationale is straightforward: models trained on similar distributions exhibit a degree of inherent alignment. AEs that capitalize on this alignment between the surrogate and black-box models are more likely to transfer. However, the persisting discrepancies [50] between models, along with the overfitting of gradient-based optimization, sharply misalign the surrogate model and black-box models on AEs. In Fig. 4, we illustrate the average alignment metric between the surrogate and black-box models (defined subsequently in Eq. (6)) across the attack iterations. It becomes apparent that direct utilization of the surrogate model’s gradients can lead to substantial misalignment with black-box models, culminating in poor transferability. Although transformations can mitigate this misalignment, it prompts the question: how can one blindly assess the efficacy of these basic transformations in facilitating this mitigation?
@@ 3.2 Assessing Effectiveness via Surrogate Self-Alignment
Figure 4: Alignment between the surrogate model and black-box models on targeted AEs, which reflects the black-box transferability resulted by different transformations (see Table IV). The vanilla gradient update, i.e., TMI, quickly misaligns the surrogate from black-box models and leads to poor transferability.
Our solution leverages what we term surrogate self-alignment induced by basic transformations. To simplify our analysis, we first examine the similarity between representations of individual samples. An illustration is depicted in Fig. 5. Given the inherent alignment between models $f$ and $g$, we can reasonably postulate that the cosine similarity of their representations, $\text{CosSim}(\bm{\phi}(\bm{x}),\bm{\psi}(\bm{x}))$, is significantly greater than zero. More informally, we may suggest that $\bm{\phi}(\bm{x})\approx\bm{\psi}(\bm{x})$. For an AE $\bm{x}^{adv}$, crafted using a weakly-transferable algorithm like TMI that has a limited impact on the black-box model $g$, we hypothesize that $\bm{\psi}(\bm{x}^{adv})\approx\bm{\psi}(\bm{x})\approx\bm{\phi}(\bm{x})$, with all three representations being markedly different from $\bm{\phi}(\bm{x}^{adv})$. This difference becomes more pronounced with each iteration of the attack. Image transformations can enhance attack effectiveness by harnessing the mutual alignment between $f$ and $g$, with the aim of making $\bm{\psi}(\bm{x}^{adv})$ closer to $\bm{\phi}(\mathcal{T}(\bm{x}^{adv}))$. In line with the principles of black-box attacks, we suggest an extension of this strategy such that $\bm{\phi}(\mathcal{T}(\bm{x}^{adv}))\approx\bm{\phi}(\bm{x})$ (self-alignment).
Figure 5: The colored numbers indicate the cosine similarity between representations generated by the surrogate model (ResNet-50) and the black-box model (ConvNeXt), where the AE $\bm{x}^{adv}$ is created based on the TMI algorithm. On the clean sample, these two models exhibit a high degree of alignment. However, AEs with weak transferability have limited impact on the black-box model, which suggests that we can introduce an image transformation $\mathcal{T}$ to make $\bm{\phi}(\mathcal{T}(\bm{x}^{adv}))$ more closely approximate $\bm{\phi}(\bm{x})$ (self-alignment), thereby enhancing the approximation to $\bm{\psi}(\bm{x}^{adv})$ (alignment) and improving transferability.
We then expand the concept of similarity between individual representations to encompass the mutual alignment across the entire dataset and multiple black-box models. Let $\Phi=\{\bm{\phi}(\bm{x}_{i})\}_{i=1}^{b}$, $\Phi_{\widehat{\mathcal{T}}_{s}}=\{\bm{\phi}(\widehat{\mathcal{T}}_{s}(\bm{x}_{i}))\}_{i=1}^{b}$, and $\Psi=\{\bm{\psi}(\bm{x}_{i})\}_{i=1}^{b}$ denote the sets of representations produced by models $f$, $f\circ\widehat{\mathcal{T}}_{s}$, and $g$ for a set of samples $\mathcal{X}=\{\bm{x}_{i}\}_{i=1}^{b}$, respectively. And $s$ represents the intensity level of a basic transformation $\widehat{\mathcal{T}}$. The superscript $(\cdot)^{adv}$ indicates the corresponding representations for the adversarial counterparts of $\mathcal{X}$, which are crafted using the TMI algorithm and model $f$. To quantify the mutual alignment between two sets of representations, such as $\Phi$ and $\Psi$, we measure the overlap of their respective $k$-nearest neighbors [51] throughout the dataset. This is formalized as:
$\operatorname{Alignment}\left(\Phi,\Psi\right)=$ (6)
@@
$\mathbb{E}_{g\in\mathcal{G},(\bm{\phi},\bm{\psi})\in(\Phi,\Psi)}$ $\frac{1}{k}\left|d_{\text{knn}}(\bm{\phi},\Phi\setminus\bm{\phi})\cap d_{\text{knn}}(\bm{\psi},\Psi\setminus\bm{\psi})\right|,$
@@
where $d_{\text{knn}}$ returns the set of indices corresponding to its $k$-nearest neighbors, with $k$ set to 100 for a dataset comprising $b=1,000$ images.
@@
TABLE I: Mutual alignment between surrogate model and multiple black-box models.
@@
Figure 6: The metrics of self-alignment and alignment exhibit a high degree of similarity. Notably, the scaling transformation demonstrates distinctive effectiveness, simultaneously inducing high alignment strength and a large effective range of transformation intensities.
Table I presents the alignment metrics between models for various combinations. The results are consistent with those illustrated in Fig. 5, demonstrating that the alignment scores for $\operatorname{Alignment}(\Phi,\Psi)$ and $\operatorname{Alignment}(\Psi,\Psi^{adv})$ are not only comparable but also significantly exceed those of the other two scenarios. This substantiates the efficacy of our motivation, leveraging the self-alignment, $\operatorname{Alignment}(\Phi^{adv}_{\widehat{\mathcal{T}}_{s}},\Phi)$, as a proxy for desired alignment, $\operatorname{Alignment}(\Phi^{adv}_{\widehat{\mathcal{T}}_{s}},\Psi^{adv})$, to blindly gauge the impact of different transformations on the transferability of attacks.
Fig. 6(a) delineates the self-alignment metric for the basic transformations under study at varying intensities. Notably, scaling not only yields a significantly enhanced surrogate self-alignment compared to other geometric transformations but also demonstrates a wider effective intensity range, highlighting its potential as a key transformation for our strategy. To corroborate these preliminary assessments and findings, we conduct comparative experiments among these basic transformations in Section 4.2.2.
Effectiveness of the surrogate self-alignment metric. We further validate the effectiveness of surrogate self-alignment in blindly estimating the real alignment towards unknown black-box alignment. Please see Figs. Fig. 6(a) and Fig. 6(b), these two metrics closely mirror each other, reflecting both the individual transformation trends at distinct intensities and the comparative magnitudes across transformations, thereby corroborating our analysis.
@@ 3.3 Revealing Correlation via Self-Transferability
We further investigate the correlations among different basic transformations, which are also achieved without utilizing the target model or its training data. Specifically, we first generate targeted AEs using diverse complex transformation methods (as described in Sec. 2.4). Then, we perform PCC tests on their self-transferability when attacking the surrogate model combined with various basic transformations. This allows us to obtain the correlation relationships between the feature diversity induced by different basic transformations. The underlying rationale is to treat basic transformations as a defensive layer [52] of the surrogate model and use their empirical expectations under different intensities as an estimation of black-box models. If an AE exhibits strong transferability against black-box models, it should also exhibit strong self-transferability against the surrogate model under basic transformation defenses. Furthermore, if the self-transferability against two basic transformations shows a strong correlation, it indicates redundancy between them. Through this analysis, we aim to provide a foundation for the combination and redundancy reduction of our method design. The self-transferability is defined as the mean efficacy of targeted AEs across varying transformation intensities:
@@
The results of PCC tests for all studied basic transformations are detailed in Fig. 7. It reveals that there are strong internal correlations within geometric transformations and within color transformations, but the correlation between these two categories is relatively weak. Given this, it is more strategic to target a few highly effective basic transformations to optimize overall performance, rather than indiscriminately combining overlapping operations. For example, considering the significant utility of the scaling transformation we previously reported, it may be more advantageous to combine it with color transformations rather than additional geometric transformations to achieve a synergistic enhancement. And the proposed self-transferability measure can assist us in conducting clear and rigorous analyses under strict black-box conditions, see Sec. 3.4.3 for details.
@@
Effectiveness of the self-transferability measure. Here we validate whether the proposed self-transferability measure, i.e., Eq. (7), serves as a high-quality estimation of the black-box transferability, i.e., Eq. (5). In Fig. 8, we present scatter plots for rotation and scaling transformations, revealing notably tighter correlations in the cross-method scenario compared to previous metrics (as in Fig. 3). The self-transferability of AEs against these basic transformations is observed as an effective reflection of the trends observed in their black-box transferability, applicable to both cross-method and intra-method contexts. The PCCs for all studied basic transformations are detailed in Fig. 9, from which we can discern the following:
@@
    The AEs’ self-transferability against most geometric transformations demonstrates a consistently high-quality linear correlation with their black-box transferability. For example, attacking features at varying frequency-domain distortions (which often induce color shifts) [46], elastic deformations [13], and random block-wise transformations [11]—despite their seemingly disparate natures—uniformly impact the exploitation of features across scales. This correlation persists even with more complex transformations and a range of geometric ones, such as rotation, shear, and translation.
@@
    The self-transferability to color transformations shows a weak correlation with black-box transferability in cross-method comparisons, but a robust correlation in the intra-method cases of SSA and SIA. This can be ascribed to the fact that both transformations pertain to the manipulation of image color.
Figure 9: PCCs calculated between the black-box transferability against 14 black-box models and the self-transferability against various basic transformations, based on AEs obtained from 12 existing attacks.
The above results demonstrate that the self-transferability against basic geometric transformations, such as rotation and scaling, can serve as an excellent estimator for black-box transferability. This makes it possible to design transformation methods and perform parameter tuning under strict black-box conditions. We leave that for future study.
3.${}^{\textbf{4}}$
Drawing on the insights and findings from our research, we introduce a novel transformation methodology focused on scaling, named Strong, Self-transferable, Fast, and Simple Scale Transformation (S4ST), designed to enhance data-free TTAs. Fig. 10 illustrates the schematic overview of S4ST, with detailed explanations to follow.
Figure 10: The proposed S4ST encompasses the Base operation, which is based on scaling, complemented by two orthogonal enhancements—Aug and Block.
3.4.1 Overview
The essence of the S4ST framework is founded on the basic scaling transformation, informed by our preceding analysis, and seeks synergistic benefits with complementary basic operations. In alignment with recent advancements, our framework also integrates strategies for image block manipulation. To accommodate block-wise scaling and alleviate the computational demands associated with upscaled inputs, we have refined the scaling transformation to preserve dimensional consistency. As a result, the S4ST framework is segmented into three distinct components: S4ST-Base, S4ST-Aug, and S4ST-Block, each reflecting one of the three primary motivations identified.

### v3

@@ 2.3 Limitations of Existing Design Consensuses
As shown in Fig. 6, intra-method analyses across hyperparameter configurations often reveal positive correlations, suggesting these metrics guide parameter scaling within specific transformations. However, severe inter-method confusion complicates designing novel transformations from scratch. This lack of reliable cross-method guidance underscores the need for a fundamental analysis, motivating our proposed self-alignment and self-transferability measures.
@@ 3 The Proposed Method
This section delineates our proposed method. We introduce two strategies to gauge the per-transformation effectiveness and cross-transformation correlations among 12 basic transformations, culminating in the powerful yet efficient S4ST for data-free TTAs.
@@ 3.1 Motivation
We aim to develop robust blind estimators for designing data-free TTA transformations, answering two core questions: 1) which transformations are most effective? and 2) how to determine their synergistic effects? Our goal is to outperform data-reliant methods in efficacy while maintaining flexibility for any (sample, label) pair without extra data or training. By adhering to strict black-box constraints, these estimators ensure applicability to broader data beyond natural images. We begin by analyzing individual basic transformations before exploring their combinations (details and intensities $\mathcal{S}$ are in the supplementary).
3.2 Assessing Effectiveness via Self-Alignment
Evaluating basic transformations’ effectiveness in enhancing targeted transferability is challenging due to the inaccessibility of black-box models and the complex dynamics of randomized parameters across hundreds of attack iterations. To address this, we propose leveraging the AE overfitting phenomenon itself as a proxy shortcut.
@@ 3.2 Assessing Effectiveness via Self-Alignment
Figure 7: Overfitting and alignment metrics between the surrogate model and black-box models on targeted AEs, which reflects the black-box transferability resulting from different transformations (see Fig. 3). The vanilla gradient update, i.e., TMI, quickly misaligns the surrogate from black-box models and leads to poor transferability.
3.2.1 Feature-space Model Alignment as Proxy
During vanilla attacks (e.g., TMI), the surrogate’s target-label confidence rapidly saturates ($f(\bm{x}_{\mathcal{T}}^{adv},t)\rightarrow 1$), causing gradient vanishing and severe overfitting (Fig. 7a), formalized as $\mathbb{E}_{\bm{x}\in\mathcal{X},g\in\mathcal{G}}\left(f(\mathcal{T}(\bm{x}_{\mathcal{T}}^{adv}),t)-g(\bm{x}_{\mathcal{T}}^{adv},t)\right)$. Conversely, applying transformations mitigates this.
Why are transformations effective? Just as the best defensive transformations restore predictions toward clean samples [89], the best surrogate transformations during attacks recalibrate the surrogate’s feature extraction to resemble that of clean samples. Beyond alleviating confidence saturation, this crucially realigns feature representations between surrogate and black-box models. This is grounded in representation convergence: well-trained models in the same domain share similar feature spaces due to common data manifolds [1, 33, 97]. AEs that preserve this inherent alignment transfer better, whereas gradient descent for a specific target disrupts it. Transformations thus help recalibrate the surrogate’s features to exploit the original alignment with black-box models.
While gradient alignment [98], $\mathbb{E}_{\bm{x}\in\mathcal{X}}\text{CosSim}\left(\nabla_{\bm{x}}\mathcal{L}_{f},\nabla_{\bm{x}}\mathcal{L}_{g}\right)$, directly measures transferability (Eq. (3)), we adopt feature-space alignment, $\mathbb{E}_{\bm{x}\in\mathcal{X}}\text{CosSim}\left(\bm{\phi}(\bm{x}),\bm{\psi}(\bm{x})\right)$, as a relaxed, label-agnostic proxy suitable for black-box settings with disparate label sets. To capture local geometry explicitly, we utilize a $k$-nearest neighbors overlap measure [33]
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\operatorname{Alignment}\left(\Phi,\Psi\right)=}$ (5)
@@ 3.2.1 Feature-space Model Alignment as Proxy
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{E}_{g\in\mathcal{G},(\bm{\phi},\bm{\psi})\in(\Phi,\Psi)}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{k}\left|d_{\text{knn}}(\bm{\phi},\Phi\setminus\bm{\phi})\cap d_{\text{knn}}(\bm{\psi},\Psi\setminus\bm{\psi})\right|,}$
@@
where $\Phi=\{\bm{\phi}(\bm{x}_{i})\}_{i=1}^{b}$ and $\Psi=\{\bm{\psi}(\bm{x}_{i})\}_{i=1}^{b}$ denote features from $f$ and $g$ for dataset $\mathcal{X}$. Function $d_{\text{knn}}$ returns the indices of the $k$-nearest neighbors, with $k=100$ for $b=1,000$ images.
Figure 8: The colored numbers indicate the cosine similarity between features generated by the surrogate model (ResNet-50) and the black-box model (ConvNeXt), where the AE $\bm{x}^{adv}$ is created based on the TMI algorithm. On the clean sample, these two models exhibit a high degree of alignment. However, an AE with weak transferability has limited impact on the black-box model, which suggests that we can introduce an image transformation $\mathcal{T}$ to make $\bm{\phi}(\mathcal{T}(\bm{x}^{adv}))$ more closely approximate $\bm{\phi}(\bm{x})$ (self-alignment), thereby enhancing the approximation towards $\bm{\psi}(\bm{x}^{adv})$ (alignment).
We validate this proxy by probing inter-model alignment $\operatorname{Alignment}(\Psi^{adv}_{\mathcal{T}},\Phi^{adv}_{\mathcal{T}})$ across attack iterations. Here, $(\cdot)^{adv}_{\mathcal{T}}$ denotes features of adversarial counterparts crafted via $\mathcal{T}$-TMI on surrogate $f$. As Fig. 7(b) shows, vanilla surrogate gradients cause rapid misalignment and poor transferability. Conversely, effective transformations maintain alignment. Crucially, this alignment metric reliably predicts the actual attack transferability (Fig. 3), validating its utility as an assessment proxy.
3.2.2 From Black-box Alignment to Self-Alignment
To circumvent dynamic probing complexities under black-box constraints, we propose self-alignment. This measures how transformations restore single-step feature consistency on overfitted AEs within the surrogate, revealing their efficacy in mitigating overfitting.
Illustrated in Fig. 8, sample-level feature similarity intuitively justifies this. Given inherent alignment $\bm{\phi}(\bm{x})\approx\bm{\psi}(\bm{x})$, a weakly-transferable AE $\bm{x}^{adv}$ (e.g., via TMI) barely impacts the black-box model $g$. Thus, $\bm{\psi}(\bm{x}^{adv})\approx\bm{\psi}(\bm{x})\approx\bm{\phi}(\bm{x})$, all diverging markedly from the overfitted surrogate feature $\bm{\phi}(\bm{x}^{adv})$ as attack iterations progress. To enhance transferability by bringing $\bm{\psi}(\bm{x}^{adv})$ closer to $\bm{\phi}(\mathcal{T}(\bm{x}^{adv}))$, we extend this strategy under strict black-box principles to pursue self-alignment: $\bm{\phi}(\mathcal{T}(\bm{x}^{adv}))\approx\bm{\phi}(\bm{x})$.
We validate this condition at the dataset level across multiple black-box models (Table I, where subscript-free sets denote vanilla TMI features). The high and comparable scores of $\operatorname{Alignment}(\Phi,\Psi)$ and $\operatorname{Alignment}(\Psi,\Psi^{adv})$ confirm our analysis. Consequently, surrogate self-alignment on severely overfitted AEs, $\operatorname{Alignment}(\Phi^{adv}_{\mathcal{T}},\Phi)$, emerges as a viable proxy for the desired black-box alignment, $\operatorname{Alignment}(\Phi^{adv}_{\mathcal{T}},\Psi^{adv})$.
@@ 3.2.2 From Black-box Alignment to Self-Alignment
TABLE I: Alignment between surrogate model and multiple black-box models.
@@
Figure 9: The metrics of self-alignment and black-box alignment exhibit a high degree of similarity. Notably, the scaling transformation demonstrates distinctive effectiveness, simultaneously inducing high alignment strength and a large effective range of transformation intensities.
Assessing effectiveness via self-alignment. Fig. 32(a) plots $\operatorname{Alignment}(\Phi^{adv}_{\widehat{\mathcal{T}}_{s}},\Phi)$ for basic transformations $\widehat{\mathcal{T}}$ across intensities $s$. While geometric transformations (rotation, scaling, perspective) promote self-alignment, scaling uniquely achieves significantly higher self-alignment over a broader effective intensit y range, marking it as a key transformation. Though self-alignment is a single-step estimation on overfitted AEs, full-attack experiments in Section 4.2.2 consistently corroborate scaling’s superior transferability.
Effectiveness of the self-alignment measure. Figs. 32(a) and 32(b) confirm that self-alignment accurately mirrors black-box alignment towards unknown models, capturing both intensity trends and relative transformation magnitudes. This reliability is rooted in representation convergence, well-trained domain models develop aligned feature spaces regardless of architecture or training distributions. Extensive cross-model evaluations (see supplementary) substantiate this core assumption, providing strong empirical support for self-alignment as a robust proxy.
@@ 3.3 Revealing Correlation via Self-Transferability
To investigate cross-transformation correlations under strict black-box constraints, merely measuring feature distribution similarities of overfitted AEs (akin to self-alignment) is insufficient, as it fails to capture the full dynamics of AE optimization.
Thus, we propose self-transferability for a more robust black-box estimation. We generate targeted AEs using diverse complex transformations (see Section 2.3 and supplementary) and compute their PCCs when attacking the surrogate defended by basic transformations. Treating these basic transformations as defensive layers [38, 89], their empirical expectations across intensities are then a proxy of black-box transferability:
$\textit{ST}_{\widehat{\mathcal{T}}}=\mathbb{E}_{\bm{x}\in\mathcal{X},s\in\mathcal{S}}\left(f(\widehat{\mathcal{T}}_{s}(\bm{x}_{\mathcal{T}}^{adv}),t)-f(\widehat{\mathcal{T}}_{s}(\bm{x}),t)\right).$ (6)
Highly correlated self-transferability between two basic transformations indicates redundancy, providing a foundation for synergistic combinations.
@@
Revealing correlation via self-transferability. Fig. 10 details the PCC tests (calculation details in supplementary). It reveals strong intra-category correlations within geometric and color transformations, but weak inter-category correlations. Consequently, it is more strategic to combine highly effective, distinct transformations, such as pairing the potent scaling transformation with color adjustments, rather than indiscriminately stacking redundant geometric operations. This measure rigorously guides our augmentation design (Section 3.4.2).
@@
Effectiveness of the self-transferability measure. We validate self-transferability (Eq. (8)) as a high-quality estimator for black-box transferability (Eq. (4)). Scatter plots for rotation and scaling (Fig. 11) reveal notably tighter cross-method correlations than prior metrics (Fig. 6), effectively reflecting black-box trends. Analyzing the PCCs across all basic transformations (Fig. 12) yields key findings:
@@
    Self-transferability against most geometric transformations demonstrates consistently high-quality linear correlations with black-box transferability. Notably, seemingly disparate attacks, like frequency-domain distortions [47], elastic deformations [41], and random block-wise transformations [78], uniformly impact multi-scale feature exploitation, maintaining strong correlations across geometric transformations like rotation, shear, and translation.
@@
    Self-transferability against color transformations exhibits weak cross-method correlations but robust intra-method correlations (e.g., SSA and SIA, both involving image color manipulation).
Figure 12: PCCs calculated between the black-box transferability against 14 black-box models and the self-transferability against various basic transformations, based on AEs obtained from 12 existing attacks under a total of 40 hyperparameter configurations.
These results confirm that self-transferability against basic geometric transformations (e.g., rotation, scaling) acts as an excellent blind estimator, enabling strict data-free method design and parameter tuning. Note that to strictly preserve the black-box setting, this specific analytical finding is excluded from our main attack experiments.
3.4 $^{\textbf{4}}S$T
Drawing on our analytical findings, we introduce Strong, Self-transferable, Fast, and Simple Scale Transformation (S4ST), a novel scaling-centered transformation designed to enhance data-free TTAs. Fig. 13 illustrates its schematic overview. S4ST is rooted in an improved scaling transformation and synergized with complementary operations and block-wise manipulation strategies, comprising three distinct components: S4ST-Base, S4ST-Aug, and S4ST-Block.
Figure 13: The proposed S4ST encompasses the Base operation, which is based on scaling, complemented by two enhancements—Aug and Block.